# INSIGHT: End-to-End Neuro-Symbolic Visual Reinforcement Learning with Language Explanations

### v1

@@ Abstract
Neuro-symbolic reinforcement learning (NS-RL) has emerged as a promising paradigm for explainable decision-making, characterized by the interpretability of symbolic policies. For tasks with visual observations, NS-RL entails structured representations for states, but previous algorithms are unable to refine the structured states with reward signals due to a lack of efficiency. Accessibility is also an issue, as extensive domain knowledge is required to interpret current symbolic policies. In this paper, we present a framework that is capable of learning structured states and symbolic policies simultaneously, whose key idea is to overcome the efficiency bottleneck by distilling vision foundation models into a scalable perception module. Moreover, we design a pipeline that uses large language models to generate concise and readable language explanations for policies and decisions. In experiments on nine Atari tasks, our approach demonstrates substantial performance gains over existing NSRL methods. We also showcase explanations for policies and decisions.
@@ 1 Introduction
Recent years have witnessed remarkable progress in the field of deep reinforcement learning (RL) (Agarwal et al., 2021; Wurman et al., 2022; Degrave et al., 2022). However, deep RL still faces limitations in sensitive domains due to its opaque nature (Milani et al., 2022). Neuro-symbolic reinforcement learning (NS-RL) is promising for overcoming the limitations (Verma et al., 2018; Coppens et al., 2019). In its simplest form, NS-RL represents policies with expressions over quantities related to objects, such as polynomials of object locations. The interpretability of an agent’s decision-making process is thus guaranteed by the clear semantics and concise formulation of the policies.
@@
For tasks with visual observations, NS-RL entrails structured state representations consisting of object locations, which means to map input pixels to objects. Zheng et al. (2022) proposed to extract the structured states from the Spatially Parallel Attention and Component Extraction (SPACE) model (Lin et al., 2020) trained on task-specific data. However, the state representations are fixed during policy learning due to the computational overhead of SPACE. That being said, the structured states are not refined with reward signals, which can lead to a significant loss in task performance.
@@
Meanwhile, symbolic policies can be intricate for general users, albeit being intrinsically interpretable. For instance, to interpret logical policies (Delfosse et al., 2023a) one has to be familiar with first-order logic, and for programmatic policies (Qiu & Zhu, 2022) one has to learn the corresponding grammars. From a practical point of view, it is imperative to attain accessibility to the general public in order to gain their trust. Nevertheless, there is a lack effort in the literature of NS-RL to make policy interpretations for non-expert users.
@@
In this paper, we present a framework called INSIGHT (INSIGHT) to address these issues. As illustrated in Fig. 1, INSIGHT is able to learn the object coordinates and symbolic policies simultaneously, and it can explain learned policies and specific decisions in natural language. The key idea for learning structured states is to distill vision foundation models into a scalable perception module that overcomes the efficiency drawback of previous approaches. We use the equation learner (EQL) (Sahoo et al., 2018) for learning symbolic policies from object coordinates, which allows both the perception module and the symbolic policy to be trained with rewards.
Moreover, we develop a pipeline for explaining policies and decisions in natural language using large language models (LLMs), which reduces the cognitive load on users to understand the construction of symbolic policies learned by INSIGHT. Through a step called concept grounding, we help LLMs to associate general concepts such as objects and motions with variables in the symbolic policies. Then, as illustrated in Fig. 1, LLMs are instruct to interpret the general patterns of decision-making based on the expressions of symbolic policies and specific decision based on specific values of object coordinates.
@@
We verify the efficacy of INSIGHT and its model design with extensive experiments on nine Atari games. INSIGHT outperforms all existing approaches for NS-RL (CGP (Wilson et al., 2018), Diffses (Yuan et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2023a)). We show that its empirical performance can be attributed to the refinement of structured states with reward information and our approach for learning symbolic policies. We also present examples of policy interpretations and decision explanations. In summary, our contributions are three-fold.
@@
    We demonstrate the efficacy of INSIGHT on nine Atari games and showcase its policy explanations.
Figure 1: For tasks with visual input, INSIGHT can simultaneously learn the coordinates of objects in observations and coordinate-based symbolic policies simultaneously, and it can interpret learned policies and specific decisions in natural language.
@@ 2 Related work
Approaches for explainable RL can be categorized into post hoc and intrinsic methods. The former generates explanations using predefined templates (Ehsan et al., 2018; Wang et al., 2019; Hayes & Shah, 2017) or saliency maps (Greydanus et al., 2018) but can be subjective and unreliable (Dazeley et al., 2023). The latter relies on interpretable function approximators, such as decision trees (Topin et al., 2021; Zhang et al., 2021), for better transparency, yet they often suffer from limited performance due to non-differentiability (Dazeley et al., 2023).
Approaches for NS-RL are intrinsically interpretable due to the construction of symbolic policies. Most existing approaches focus on tasks with low-dimensional observations that have clear semantics (Verma et al., 2018; Coppens et al., 2019; Verma et al., 2019; Landajuela et al., 2021). For tasks with visual input, they either rely on ground truth (Delfosse et al., 2023a) or human-defined primitives (Wilson et al., 2018; Lyu et al., 2019) for state representations. As an exception, Zheng et al. (2022) proposed to extract structured states from pre-trained SPACE models, yet the representations are not refined with reward signals. To our knowledge, INSIGHT is the first framework that learns structured states from both visual and reward signals.
Meanwhile, there is a growing interest in using LLMs to explain machine learning models. For instance, Kroeger et al. (2023) leveraged LLMs to rank features by their importance, and Tennenholtz et al. (2023) used LLMs to explain latent embeddings. INSIGHT is related to the approaches that generate explanations based on input-output pairs (Singh et al., 2023; Bills et al., 2023) or decision trees distilled from the neural models (Zhang et al., 2023). However, INSIGHT is able to prompt LLMs with the actual symbolic policies due to their simplicity and generate more informed explanations.
@@ 3.1 Visual Perception
Frame-Symbol Dataset  The perception module needs to extract information about objects from input images. Given the notorious sample efficiency of online RL, approaches that use image reconstruction objectives (Zheng et al., 2022; Yoon et al., 2023) are prohibitively expansive. We claim that, for NSRL, compared to reproducing every detail in visual observations, it suffices to only recognize the positions of objects. To this end, we consider harvesting the segmentation and tracking ability of vision foundation models. For each task, we first roll out about 10,000 observations using pre-trained neural agents, and then we extract the bounding boxes of objects in these observations using the FastSAM segmentation model (Zhao et al., 2023) and the DeAot model (Yang & Yang, 2022). Using the bounding boxes, we compute the coordinates of objects (i.e., their centers) and their shape, normalize them to $[0,1]$, and pair them with the corresponding images to form a frame-symbol dataset $\mathcal{D}_{\text{symbol}}$. Details of the dataset generation are provided in Sec. A.1. By fitting to this dataset, the perception module extract the structured states efficiently, a critical condition of refining them during policy learning.
Learning Object Coordinates  We now present the details of learning from $\mathcal{D}_{\text{symbol}}$. $\mathcal{D}_{\text{symbol}}$ contains information of the existence, the shape, and the coordinates of objects. To make full use of them, INSIGHT employs a multi-task formulation for the perception module $\omega_{\text{perception}}$, which is illustrated in Fig. 2. An encoder, parameterized by convolutional neural networks (CNNs) and FCNs, is responsible for encoding information about objets and rewards into hidden representations. Three FCNs use the hidden representations to predict the existence, the shape, and the coordinates of objects, which are explained below.
@@
$\mathcal{L}_{\mathrm{exist}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}\bar{\eta}_{ij}\Big{[}$ $c_{ij}(1-p_{ij})^{\psi}\log(p_{ij})$ (1)
@@
where $\psi$ is the modulating factor of the focal loss (Lin et al., 2017). Details for computing $\bar{\eta}_{ij}$ is provided in Sec. A.2.
@@
$\mathcal{L}_{\mathrm{coor}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}c_{ij}\left[\left|x_{i,2j}-\hat{x}_{i,2j}\right|+\left|x_{i,2j+1}-\hat{x}_{i,2j+1}\right|\right].$ (2)
Note that the predicted coordinates are clipped to $[0,1]$ when being used for policy learning, as they are supposed to be the coordinates of objects. For shape prediction, the network is trained to predict the width and height of object bounding boxes. The loss, $\mathcal{L}_{\mathrm{size}}$, takes a similar form as $\mathcal{L}_{\mathrm{coor}}$, except that the prediction targets are replaced with the width and height of objects. Overall, the objective for learning from the frame-symbol dataset is:
@@ 3.2 Learning Symbolic Policies
where $f_{1}$, $f_{2}$, …, $f_{d^{\prime}_{k}}\in\mathcal{F}$ are scalar functions. $f_{1},f_{2},\dots,f_{u}$ are unary functions, and the remaining are binary functions. EQL networks are often much smaller than FCNs to enhance interpretability. It is the flexibility of activation functions that enhances expressiveness. In this work, $\mathcal{F}$ includes the square function, the cubic function, constants, the identity function, multiplication, and addition.
@@
Remark  By using a clipping objective, the PPO algorithm supports reusing the same online samples for multiple iterations to improve sample efficiency However, $\mathcal{L}_{\text{cnn}}$ and $\mathcal{L}_{\text{ng}}$ are not defined as clipping objectives. In preliminary experiments, we found that optimizing $\mathcal{L}_{\text{cnn}}$ and $\mathcal{L}_{\text{ng}}$ for multiple iterations led to inferior performance, possibly due to drastic changes imposed on the encoder of the perception module $\omega_{\text{perception}}$. In consequence, we optimize $\mathcal{L}$ in the last iteration and $\mathcal{L}_{\text{ppo}}$ in all other iterations in Alg. 1.
Table 1: Performance of the proposed INSIGHT, Neural (a neural agent that has the same network architecture as INSIGHT), and existing NS-RL approaches (CGP (Wilson et al., 2018), Diffses (Zheng et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2023a)) on nine Atari tasks. INSIGHT outperforms all NS-RL baselines and matches the performance of Neural for all nine tasks, which means it overcomes the performance drawback of previous NS-RL approaches.
@@
BeamRider $3828.1\pm 261.1$ $\textbf{3868.1}\pm 204.2$ $1614.4\pm 79.8$ $1341.6\pm 21$ / $354.6\pm 20.3$ / 5775
Enduro $843.7\pm 213.8$ $676.9\pm 730.4$ $\textbf{933.9}\pm 4.2$ $2\pm 0$ / $41.1\pm 15.5$ $2.4\pm 2.1$ 309.6
Qbert $16978.6\pm 1936.1$ $\textbf{17879.2}\pm 1857.1$ $13269.2\pm 1024.4$ $770\pm 94$ / $558.3\pm 41.5$ / 13455
@@
    Since DSP cannot handle visual observations, we used pre-trained peception module of INSIGHT to extract object coordinates and used its released code for policy learning.
@@
    We used the results reported by Delfosse et al. (2023a) for Freeway and obtained results for Pong, Enduro, SpaceInvaders, Seaquest, and Breakout using its released code. Due to the absence of predefined templates for all tasks within the codebase, we tailored the template originally designed for Asterix to fit the unique action spaces of the additional tasks. We were unable to obtain results for BeamRider, Qbert, and MsPacman since the ground-truth object locations are not available for them.
@@ 3.3 Explaining Policies and Decisions
The pipeline for generating explanations starts with a step called concept grounding, which is followed by separate prompts for policy interpretations and decision explanations. Full prompts and details are provided in Appx. D.
Suppose a user who is familiar with a task wants to interpret a learned symbolic policy. What can be tedious for the user is to associate concepts of the task, such as the goal and the influence of actions, with the construction of the symbolic policy. For example, he or she has to find out which element in the coordinate vector corresponds to the location of a certain object. We refer the process of establishing such a correspondence as concept grounding. Note that concept grounding is in fact not necessary for understanding explanations for policies and decisions. For example, only knowledge of Pong is required to understand the conclusion part in the right part of Fig. 5, which are explanations for choosing action up at a particular state. This observation inspires us to use LLMs to ground concepts and generate explanations, thereby reducing the user’s intellectual burden.
@@
As shown in the upper left of Fig. 5, learned policies are formulated as a set of equations involving input variables (i.e. object coordinates), intermediate variables (e.g., $t_{1}$ in Fig. 5), and action logits. Despite their simplicity, prompting LLMs with these equations only yields vacuous responses such as "the policy is complex, involving non-linear combinations of input variables". Therefore, we apply the chain-of-thought principle and proceed in three steps: analyze the mapping from input variables to intermediate variables, analyze the mapping from intermediate variables to action logits, and finally summarize the findings.

### v2

@@ Abstract
Neuro-symbolic reinforcement learning (NS-RL) has emerged as a promising paradigm for explainable decision-making, characterized by the interpretability of symbolic policies. NS-RL entails structured state representations for tasks with visual observations, but previous methods are unable to refine the structured states with rewards due to a lack of efficiency. Accessibility also remains to be an issue, as extensive domain knowledge is required to interpret symbolic policies. In this paper, we present a framework for learning structured states and symbolic policies jointly, whose key idea is to distill vision foundation models into a scalable perception module and refine it during policy learning. Moreover, we design a pipeline to generate language explanations for policies and decisions using large language models. In experiments on nine Atari tasks, we verify the efficacy of our approach, and we also present explanations for policies and decisions.
@@ 1 Introduction
Recent years have witnessed remarkable progress of deep reinforcement learning (RL) (Agarwal et al., 2021; Wurman et al., 2022; Degrave et al., 2022). However, deep RL still faces limitations in sensitive domains due to its opaque nature (Milani et al., 2022). For example, the opacity hinders diagnosing or rectifying policies that fails to generalize, which may be caused by exploiting peripheral information (Delfosse et al., 2024). Neuro-symbolic reinforcement learning (NS-RL) is promising for overcoming this limitation (Verma et al., 2018; Coppens et al., 2019). It uses structured states and parameterizes policies with concise expressions, thereby guaranteeing interpretability via clear semantics of states and policies.
@@
For tasks with visual observations, the structured states need to be learned from pixels, which involves identifying objects from images. Zheng et al. (2022) proposed to use the Spatially Parallel Attention and Component Extraction (SPACE) model (Lin et al., 2020) for this purpose, but the state representations are fixed during policy learning due to the computational overhead of SPACE. That being said, the structured states are not refined with reward signals, leading to significant performance degradation.
@@
Meanwhile, symbolic policies can be intricate for general users, albeit being intrinsically interpretable. For instance, to interpret logical policies (Delfosse et al., 2023b) one has to be familiar with first-order logic, and in the case of programmatic policies (Qiu & Zhu, 2022) one has to learn the corresponding grammars. Such a lack of accessibility can be an obstacle for agents to gain the trust of the general public. Nevertheless, there is a lack effort in the literature of NS-RL to explain policies for non-expert users.
@@
In this paper, we present a framework called INSIGHT (INSIGHT) to address these issues. As illustrated in Fig. 1, INSIGHT can learn the object coordinates and symbolic policies simultaneously and explain policies and specific decisions in natural language. The key idea of INSIGHT is to overcomes the efficiency drawback of previous methods by distilling vision foundation models into a scalable perception module. Accompanied by the equation learner (EQL) (Sahoo et al., 2018) for representing policies, both the perception module and symbolic policies can be learned from rewards in an end-to-end fashion. Moreover, we develop a pipeline for explaining policies and decisions in natural language using large language models (LLMs), which reduces the cognitive load on users to understand symbolic policies learned by INSIGHT. Through a step called concept grounding, LLMs are instructed to associate quantities in symbolic policies with their semantics for the task of interest. Then, as illustrated in Fig. 1, LLMs are prompted for explaining the decision-making patterns of a symbolic policy and specific decisions made by the policy.
@@
We verify the efficacy of INSIGHT and its model design with extensive experiments on nine Atari games. INSIGHT outperforms all existing approaches for NS-RL (CGP (Wilson et al., 2018), Diffses (Yuan et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2023b)). We show that its empirical performance can be attributed to the refinement of structured states with reward information and our approach for learning symbolic policies. We also present examples of policy interpretations and decision explanations. In summary, our contributions are three-fold.
@@
    We demonstrate the efficacy of our framework on nine Atari games and showcase language explanations.
Figure 1: For tasks with visual input, INSIGHT can simultaneously learn the coordinates of objects in observations and coordinate-based symbolic policies simultaneously, and it can interpret learned policies and specific decisions in natural language. $\text{y}_{\text{{agent,i}}}$ represents the vertical coordinate of the agent in the $i^{th}$ frame. Both policy interpretation and decision explanation are produced by entering the policy and a predefined prompt template into the LLM.
@@ 2 Related work
Approaches for explainable RL can be categorized into post hoc and intrinsic methods. The former generates explanations using predefined templates (Ehsan et al., 2018; Wang et al., 2019; Hayes & Shah, 2017) or saliency maps (Greydanus et al., 2018), which can be subjective and unreliable (Dazeley et al., 2023). The latter relies on interpretable policy classes, such decision trees (Topin et al., 2021; Zhang et al., 2021), logical expressions (Dazeley et al., 2023; Delfosse et al., 2023b, 2024), and mathematical expressions (Zheng et al., 2022; Landajuela et al., 2021), for better transparency. Meanwhile, we argue that intrinsic transparency does not guarantee accessibility. While logic-based approaches are better for characterizing relations, they can be unfriendly to users due to their prerequisite knowledge. Recently, there is a growing interest in using LLMs to explain machine learning models (Kroeger et al., 2023; Tennenholtz et al., 2023; Singh et al., 2023; Bills et al., 2023; Zhang et al., 2023). INSIGHT is the first NS-RL method that improves accessibility by using LLMs to generate explanations automatically.
Meanwhile, prior NS-RL approaches (Verma et al., 2018; Coppens et al., 2019; Verma et al., 2019; Landajuela et al., 2021) focus on tasks with low-dimensional observations that have clear semantics. For tasks with visual input, they either rely on ground truth (Delfosse et al., 2023b) or human-defined primitives (Wilson et al., 2018; Lyu et al., 2019) for state representations. As an exception, Zheng et al. (2022) proposed to extract structured states from pre-trained SPACE models, yet the representations are not refined with reward signals and lead to significant performance degradation. INSIGHT is the first NS-RL approach that learns the structured states from both visual and reward signals.
@@ 3.1 Visual Perception
Frame-Symbol Dataset  The perception module needs to extract information about objects from input images. Given the notorious sample efficiency of online RL, approaches that use image reconstruction objectives (Zheng et al., 2022; Yoon et al., 2023) are prohibitively expansive. We claim that, for NS-RL, compared to reproducing every detail in visual observations, it suffices to only recognize the positions of objects. To this end, we consider harvesting the segmentation and tracking ability of vision foundation models. For each task, we first roll out about 10,000 frames using pre-trained neural agents and then extract the object bounding boxes using the FastSAM segmentation model (Zhao et al., 2023) and the DeAot model (Yang & Yang, 2022). In specific, segmentation is repeated every ten frames to capture new objects, and the DeAot model is also capable of capturing unseen objects during tracking. Using the bounding boxes, we compute the coordinates of objects (i.e., their centers) and their width and height, normalize these quantities to $[0,1]$, and pair them with the corresponding images to form a frame-symbol dataset $\mathcal{D}_{\text{symbol}}$. Details of the dataset generation are provided in Sec. A.1. By fitting to this dataset, the perception module can learn to extract the structured states before policy learning, which can improve the sample efficiency of INSIGHT.
Learning Object Coordinates  We now present the details of learning from $\mathcal{D}_{\text{symbol}}$. $\mathcal{D}_{\text{symbol}}$ contains information of the existence, the shape, and the coordinates of objects. To make full use of them, INSIGHT employs a multi-task formulation for the perception module $\omega_{\text{perception}}$, which is illustrated in Fig. 2. An encoder, parameterized by convolutional neural networks (CNNs) and FCNs, is responsible for encoding information about objects and rewards into hidden representations. Three FCNs use the hidden representations to predict the existence, the shape, and the coordinates of objects, which are explained below.
@@
$\mathcal{L}_{\mathrm{exist}}=\sum_{j=1}^{C}\bar{\eta}_{ij}\Big{[}$ $c_{ij}(1-p_{ij})^{\psi}\log(p_{ij})$ (1)
@@
[…]
@@
$\mathcal{L}_{\mathrm{coor}}=\sum_{j=1}^{C}c_{ij}\left[\left|x_{i,2j}-\hat{x}_{i,2j}\right|+\left|x_{i,2j+1}-\hat{x}_{i,2j+1}\right|\right].$ (2)
These losses are then averaged over all frames. Note that the predicted coordinates are clipped to $[0,1]$ when being used for policy learning, as they are supposed to be the coordinates of objects. For shape prediction, the network is trained to predict the width and height of object bounding boxes. The loss, $\mathcal{L}_{\mathrm{size}}$, takes a similar form as $\mathcal{L}_{\mathrm{coor}}$, except that the prediction targets are replaced with the width and height of objects. Overall, the objective for learning from the frame-symbol dataset is:
@@ 3.2 Learning Symbolic Policies
where $f_{1}$, $f_{2}$, …, $f_{d^{\prime}_{k}}\in\mathcal{F}$ are scalar functions. $f_{1},f_{2},\dots,f_{u}$ are unary functions, and the remaining are binary functions. This flexible choice of activation functions enhances expressiveness of EQL. In this work, $\mathcal{F}$ includes the square function, the cubic function, constants, the identity function, multiplication, and addition.
@@
Remarks  The PPO algorithm reuses online samples for multiple iterations to improve sample efficiency. In preliminary experiments, we found that optimizing $\mathcal{L}_{\text{cnn}}$ and $\mathcal{L}_{\text{ng}}$ for multiple iterations led to inferior performance, possibly due to drastic changes imposed on the perception module $\omega_{\text{perception}}$. Thus, we optimize $\mathcal{L}$ only in the last iteration and $\mathcal{L}_{\text{ppo}}$ in all other iterations in Alg. 1.
The perception module $\omega_{\text{perception}}$ sacrifices some auxiliary information for the sake of efficiency, since we only include the location, heights, and width in $\mathcal{D}_{\text{symbol}}$. This is where end-to-end policy learning comes in. By refining $\omega_{\text{perception}}$ using reward signals, the hidden layers of $\omega_{\text{perception}}$ can learn to capture features that are essential for task performance but overlooked during pre-training. As shown by our results presented in Tab. 4 and Fig. 3, end-to-end policy improves both task performance and coordinate prediction.
Table 1: INSIGHT matches the performance of neural agents on all Atari tasks. Performance of the proposed INSIGHT, Neural (a neural agent that has the same network architecture as INSIGHT), and existing NS-RL approaches (CGP (Wilson et al., 2018), Diffses (Zheng et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2023b)) on nine Atari tasks. INSIGHT outperforms all NS-RL baselines and matches the performance of Neural for all nine tasks.
@@
BeamRider $3828.1\pm 261.1$ $3868.1\pm 204.2$ $1614.4\pm 79.8$ $1341.6\pm 21$ / $354.6\pm 20.3$ / 5775
Enduro $843.7\pm 213.8$ $676.9\pm 730.4$ $933.9\pm 4.2$ $2\pm 0$ / $41.1\pm 15.5$ $2.4\pm 2.1$ 309.6
Qbert $16978.6\pm 1936.1$ $17879.2\pm 1857.1$ $13269.2\pm 1024.4$ $770\pm 94$ / $558.3\pm 41.5$ / 13455
@@
    Since DSP cannot handle visual observations, we used pre-trained peception module of INSIGHT to extract object coordinates and used its released code for policy
    learning.
@@
    We used the results reported by Delfosse et al. (2023b) for Freeway and obtained results for Pong, Enduro, SpaceInvaders, Seaquest, and Breakout using its released code.
    Due to the absence of predefined templates for all tasks within the codebase, we tailored the template originally designed for Asterix to fit the unique action spaces of the
    additional tasks. We were unable to obtain results for BeamRider, Qbert, and MsPacman since the ground-truth object locations are not available for them.
@@ 3.3 Explaining Policies and Decisions
The pipeline for generating explanations starts with a step called concept grounding, which is followed by separate prompts for policy interpretations and decision explanations. Full prompts and details are provided in Appx. D. Suppose a user who is familiar with a task wants to interpret a learned symbolic policy. What can be tedious for the user is to associate concepts of the task, such as the goal and the influence of actions, with the construction of the symbolic policy. For example, he or she has to find out which element in the coordinate vector corresponds to the location of a certain object. We refer the process of establishing such a correspondence as concept grounding. Note that concept grounding is in fact not necessary for understanding explanations for policies and decisions. For example, only knowledge of Pong is required to understand the conclusion part in the right part of Fig. 5, which are explanations for choosing action up at a particular state. This observation inspires us to use LLMs to ground concepts and generate explanations, thereby reducing the user’s intellectual burden.
@@
As shown in the upper left of Fig. 5, learned policies are formulated as a set of equations involving input variables (i.e. object coordinates), intermediate variables (e.g., $t_{1}$ in Fig. 5), and action logits. Despite their simplicity, prompting LLMs with these equations only yields vacuous responses such as "the policy is complex, involving non-linear combinations of input variables". Therefore, we apply the chain-of-thought principle and proceed in three steps: analyze the mapping from input variables to intermediate variables, analyze the mapping from intermediate variables to action logits, and finally summarize the findings. Moreover, to mitigate issues such as hallucinations and uncontrollable outputs, we carefully design a set of rules for the LLM, as detailed in Tab. A16. For example, Rule 1 insists that explanations remain rooted in the policy’s expressions, and rule 2 dictates that LLMs’ policy explanations should be based on logits and action probabilities. In our preliminary experiments, these rules were able to improves the alignment between generated explanations and actual policy actions and makes outputs more precise and controllable.